Jason estimates that his car loses 12% of its value every year. The initial value is $12,000. Which best describes the graph of the function that represents the value of the car after x years?

f(x) = 12,000(0.12)x, with a horizontal asymptote of y = 0

f(x) = 12,000(1.12)x, with a vertical asymptote of x = 0

f(x) = 12,000(0.88)x, with a horizontal asymptote of y = 0

f(x) = (12,000  0.88)x, with a vertical asymptote of x = 0


1
Expert's answer
2021-12-06T16:21:54-0500
"f(x)=f(0)(1-r)^x"

"f(x)=12000(1-0.12)^x"

"f(x)=12000(0.88)^x"

"f(x)\to0\text{ as } x\to\infin"

The horizontal asymptote is "y=0."


f(x) = 12,000(0.88)x, with a horizontal asymptote of y = 0.